Find a parametric representation g(s,t) of the part of the one-sheeted hy-perboloid x^2/a^2+y^2/a^2−z^2/c^2= 1 that lies between z=−1 and z= 1
1
Expert's answer
2019-11-04T09:39:40-0500

If in the form


"{x^2 \over a^2}+{y^2\over b^2}-{z^2 \over c^2}=1"

then let "z=cv, v" as one parameter, and then, the equation can be rewritten in the following way:


"{x^2 \over a^2}+{y^2\over b^2}-{c^2v^2 \over c^2}=1""{x^2 \over a^2}+{y^2\over b^2}=1+v^2"

Let "x=a\sqrt{1+v^2}\cos u, y=b\sqrt{1+v^2}\sin u." Then

"{(a\sqrt{1+v^2}\cos u)^2 \over a^2}+{(b\sqrt{1+v^2}\sin u)^2\over b^2}=1+v^2,"

"x=a\sqrt{1+v^2}\cos u,""y=b\sqrt{1+v^2}\sin u,""z=cv,""u\in[0, 2\pi), v\in\big [-{ 1\over c} , {1 \over c} \big]."

Other parameterizations include


"x=a\cosh{v}\cos{u},"

"y=b\cosh{v}\sin{u},"

"z=c\sinh{v},"

"u\in[0, 2\pi), v\in\big [-\sinh^{-1}({ 1\over c}) , \sinh^{-1}({ 1\over c})]."